Solve for x:
X + 8 over 6 = 2 - 5x over 3

Answer by tommyt3rd(5050)

6(x +8) / 6 = 6(2-5x) / 3
x+8=2(2-5x)
x+8=4-10x
11x=-4
x=-4/11